9) What is the molarity of a solution prepared by diluting 35.8mL of 6.5 MHCI to 250.0 mL? (2pts)

please help
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 6.5 M

V1 = 35.8 mL

V2 = 250 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (6.5*35.8)/250

M2 = 0.9308 M

Answer: 0.931 M